I'm having a lot of trouble with this one:
The angle of elevation of a 15 ft ladder is 70 degrees. Find out how far the base of the ladder is from the wall.

Answer by RAY100(1637)

A rough sketch might help
.
It is a right triangle, with hypotenuse = 15, and one angle =70 deg
.
Let the distance to the wall be x
.
the trig function is adjacent / hyp or cos
.
therefore,,,,cos 70 = x/15,,,or x=15 cos 70 = 5.13
.
.
check. cos 70 = 5.13 / 15 = .3420,,,,ok
.